Water flows through a pipe of varying cylindrical cross-section. The speed of the water is 4m/s when the pipe radius 3cm. What is the speed of water through a cross-section of 6cm radius?


1
Expert's answer
2020-12-23T10:25:03-0500
"A_1v_1=A_2v_2\\\pi r_1^2v_1=\pi r_2^2v_2\\3^2(4)=6^2v_2\\v_2=1\frac{m}{s}"